OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)2 – 25? Identify the x–intercepts by graphing the quadratic equation. Answer a.x–intercepts: (2, 0) and (–8, 0) y–intercept: (0, –16) b.x–intercepts: (–2, 0) and (8, 0) y–intercept: (0, –16) c.x–intercepts: (–2, 0) and (8, 0) y–intercept: (0, –25) d.x–intercepts: (2, 0) and (–8, 0) y–intercept: (0, –25)

OpenStudy (anonymous):

First I would identify the y-int. Let x=0 y = (0 – 3)^2 – 25 y = 9-25 y = -16 so the y-int is (0, -16)

OpenStudy (anonymous):

Let y = 0 for the x-intercepts 0 = (x – 3)^2 – 25 25 = (x - 3)^2 5 = x - 3 AND -5 = x - 3 x = 8 and x = -2 The answer is b